You purchased 361 shares of stock in XYZ Company for $33 per share one year ago. You just sold your shares for $46 per share. You paid a 9% commission on each transaction. What is your rate of return for the year? Express your % answer in decimal format: 4.5% = 0.045
Answer #1

Purchase Cost = 361*33 = $11,913

Selling value (361*46)    = $16,606

Less: sales commission (16606*9% = $1494.54

Net Payment received    = $15111.46

Rate of return for the year = (Net Payment received - Purchase Cost) / Purchase Cost

= (15111.46 - 11913) / 11913

= 26.85% or 0.269